OpenStudy (anonymous):

determine the sum of the first 20 terms of the arithmetic series in which the 15th term si 107 and the terms decrease by 3. plz help, i've got a test tmrw. i need to see the procedure

OpenStudy (neer2890):

you mean to say that 14th term is 104 or 16th term is 104?

OpenStudy (anonymous):

i think the 16th term is 104

OpenStudy (anonymous):

14th is 110

OpenStudy (neer2890):

ok.considering that increasing terms decrease by 3 so, we have \[a_{15}=107\] for nth term in an arihthmetic progression, \[a _{n}=a+(n-1)d\] where a is the first term and d is the common difference. \[a _{15}=a+(15-1)(-3)\] 107=a-42 a=149

OpenStudy (neer2890):

similarly as we have to determine the sum of first 20 terms \[a _{20}=a+(20-1)(-3)\] so \[a _{20}=149-57=92\] now to find sum of terms \[S _{n}=\frac{ n }{ 2 }[a+l]\] where n is the number of terms, a is first term and l is last term.

OpenStudy (neer2890):

\[S _{20}=\frac{ 20 }{ 2 }[149+92]\] =10*241=2410

OpenStudy (neer2890):

alternatively , you can use this formula \[S _{n}=\frac{ n }{ 2 }[2a+(n-1)d]\] where n is the number of terms , a is first term and d is common difference.

OpenStudy (anonymous):

we already solved for a in the first part, so why did we solve for a again?

OpenStudy (neer2890):

\[S _{20}=\frac{ 20 }{ 2 }[2*149+(20-1)(-3)]=2410\]

OpenStudy (anonymous):

first you got 149 n then 92

OpenStudy (neer2890):

in above formula we have to find first and last terms to find the sum. so later we found 20th term of the series which is 92

OpenStudy (anonymous):

ohh...i see

OpenStudy (anonymous):

i get it, thank you!

OpenStudy (neer2890):

you are welcome.. @qwertykeyboard